if ben invests $4000 at 4% interest per year then how much additional money must he invest at 5.5% annual interest to ensure that the interest he receives each year is 4.5% of the total amount invested?
Answer by macston(5194)

.
F=amount at 5.5%
.
0.04($4000)+0.055(F)=0.045($4000+F)
$160+0.055F=$180+0.045F
0.010F=$20
F=$2000
ANSWER: He should invest $2000 at 5.5%.
.
CHECK:
0.04($4000)+0.055($2000)=0.045($4000+$2000)
$160+$110=0.045($6000)
$270=$270